OpenStudy (anonymous):

Which fraction is the greatest 3/4, 7/4, 1 1/4

OpenStudy (anonymous):

Is the last one eleven divided by four?

OpenStudy (anonymous):

No its one and 1/4. Sorry

OpenStudy (anonymous):

S'ok. Then it's 7/4

OpenStudy (anonymous):

That's what I thought. Thank you

jimthompson5910 (jim_thompson5910):

This is because 1 1/4 = 1+1/4 = 4/4+1/4 = 5/4 So 1 1/4 = 5/4

OpenStudy (anonymous):

you divide them, so 3/4 would be .75. 7/4 would be 1.75 and 1 1/4 would be 1.25 . so 7/4 would be the greatest one